Eva invested a certain amount of money at 8% interest and $1500 more than that amount at 12%. Her total yearly interest was $815. How much did she invest at each rate?
Answer by checkley79(3341)

.08X+.12(X+1,500)=815
.08X+.12X+180=815
.20X=0=815-180
.20X=635
X=635/.20
X=3,175 IS THE AMOUNT INVESTED @ 8%
3,175+1,500=4,675 IS THE AMOUNT INVESTED @ 10%.
PROOF:
.08*3,175+.12*4,675=815
254+561=815
815=815